(x+2)(x+3)=(x+7)(x-4)
Answer by mathworm7271(5)

First the left side...
(x+2)(x+3)
Use the distributive property
X^2+2x+3x+6
simplify
x^2+5x+6
Now the right side
(x+7)(x-4)
Use the distributive property
X^2+7x-4x-28
Simplify
X^2+3x-28
Put them together...
X^2+5x+6=X^2+3x-28
Subtract x^2 on both sides
5x+6=3x-28
Subtract 3x on both sides
2x+6=-28
Subtract 6 on both sides
2x=-34
Divide both sides by 2
x=-17